\begin{thebibliography}{99}

\bibitem{a:go} A. Nesip; G. Ali; \"O. Erdal, {\it On spe\-cial weakly
Ricci-sym\-metric Ken\-motsu manifolds}, Sarajevo J. Math., 3(15)(2007), no. 1, 93-97. \\ MR2327508 (2008c:53084)

\bibitem{b:tdt} T. Q. Binh; L. Tam\'assy; U. C. De; M. Tarafdar, {\it
Some remarks on almost Kenmotsu manifolds}, Math. Pannon., 13(2002), no. 1, 31-39. MR1888318 (2002k:53156)

\bibitem{d:c} D. Chinea, {\it On horizontally conformal $(\phi,\phi')$-holomorphic submersions},
Houston J. Math., 34(2008), no. 3, 721-737. MR2448378

\bibitem{c:k} M. C. Chaki, {\it On pseudo Ricci symmetric manifolds}, Bulgar. J. Phys., 15(1988), no. 6,
526-531. MR1028590 (90k:53071)

\bibitem{l:d2} L. S. Das, {\it Second order parallel tensors on $\alpha $-Sasakian manifold},
Acta Math. Acad. Paedagog. Nyh\'azi. (N.S.), 23(2007), no. 1, 65-69 (electronic). MR2322902

\bibitem{d:e} U. C. De, {\it Second order parallel tensors on $P$-Sasakian manifolds}, Publ. Math. Debrecen,
49(1996), no. 1-2, 33-37. MR1416302 (97e:53089)

\bibitem{d:m} U. C. De; A. K. Mondal, {\it On 3-dimensional normal almot contact metric manifolds satisfying certain curvature conditions},
 Commun. Korean Math. Soc., 24(2009), no. 2, 265-275.

\bibitem{d:t} U. C. De;  M. M. Tripathi, {\it Ricci tensor in 3-dimensional trans-Sasakian manifolds},
Kyungpook Math. J., 43(2003), no. 2, 247-255. MR1982228 (2004d:53049)

\bibitem{l:e} L. P. Eisenhart, {\it Symmetric tensors of the second order whose first covariant
derivatives are zero}, Trans. Amer. Math. Soc. 25 (1923), no. 2, 297-306. MR1501245

\bibitem{g:s} A. Ghosh; R. Sharma, {\it Some results on contact metric ma\-ni\-folds},
Ann. Global Anal. Geom., 15(1997), no. 6, 497-507. MR1608675 (99d:53029)

\bibitem{g:sc} A. Ghosh; R. Sharma; J. T. Cho, {\it Contact metric manifolds with $\eta $-parallel
torsion tensor}, Ann. Global Anal. Geom., 34(2008), no. 3, 287-299. MR2434858

\bibitem{j:s} S. K. Jana; A. A. Shaikh, {\it On quasi-conformally flat weakly Ricci symmetric
manifolds}, Acta Math. Hungar., 115(2007), no. 3, 197–214. MR2317216 (2008a:53042)

\bibitem{x:j} X.-q. Jia, {\it Second order parallel tensors on quasi-constant curvature manifolds},
Chinese Quart. J. Math., 17(2002), no. 2, 101-105. MR1951350 (2003h:53046)

\bibitem{j:dp}  J.-B. Jun; U. C. De; Goutam Pathak, {\it On Kenmotsu manifolds},
J. Korean Math. Soc. 42 (2005), no. 3, 435-445. MR2134708 (2006d:53044)

\bibitem{k:k} K. Kenmotsu, {\it A class of almost contact Riemannian manifolds},
T\^ohoku Math. J., 24(1972), 93-103. MR0319102 (47 \#7648)

\bibitem{k:pt} J.-S. Kim; R. Prasad; M. M. Tripathi, {\it On ge\-ne\-ra\-lized Ricci-recurrent
trans-Sasakian manifolds}, J. Korean Math. Soc. 39 (2002), no. 6, 953-961. MR1932790 (2003f:53066)

\bibitem{l:k} J. Levine; G. H. Katzin, {\it On the number of special quadratic first integrals in
affinely connected and Riemannian spaces}, Tensor (N.S.), 19(1968), 113-118. MR0224022 (36 \#7069)

\bibitem{h:l} H. Levy, {\it Symmetric tensors of the second order whose covariant derivatives vanish},
Ann. of Math., (2) 27(1925), no. 2, 91-98. MR1502714

\bibitem{z:l} Z. Li, {\it Second order parallel tensors on $P$-Sasakian manifolds with a coefficient $k$},
Soochow J. Math., 23(1997), no. 1, 97-102.  MR1436425 (97k:53044)

\bibitem{v:m} V. Mangione, {\it Harmonic maps and stability on $f$-Kenmotsu manifolds},
Internat. J. Math. Math. Sci., 2008, Art. ID 798317, 7 pp. MR2377360 (2008m:53161)

\bibitem{o:r} Z. Olszak; R. Rosca, {\it Normal locally conformal almost cosymplectic ma\-ni\-folds},
Publ. Math. Debrecen, 39(1991), no. 3-4, 315-323. MR1154263 (93c:53021)

\bibitem{o:c} C. Oniciuc, {\it Nonlinear connections on tangent bundle and harmonicity}, Ital. J. Pure Appl.
Math., 6(1999), 109-122 (2000). MR1758536 (2001e:53026)

\bibitem{n:p1} N. Papaghiuc, {\it Semi-invariant submanifolds in a Kenmotsu manifold}, Rend. Mat., (7) 3(1983),
no. 4, 607-622. MR0759118 (85i:53024)

\bibitem{n:p2} N. Papaghiuc, {\it On the geometry of leaves on a semi-invariant $\xi\sp \perp$-submanifold in a
Kenmotsu manifold}, An. Stiint. Univ. "Al. I. Cuza" Ia\c si, 38(1992), no. 1, 111-119. MR1282989 (95b:53071)

\bibitem{gh:p} Gh. Pitis, {\it Geometry of Kenmotsu manifolds}, Publishing House of Transilvania University of Bra\c sov,
Bra\c sov, 2007. MR2353263 (2008i:53117)

\bibitem{r:s1} R. Sharma, {\it Second order parallel tensor in real and complex space forms},
Internat. J. Math. Math. Sci., 12(1989), no. 4, 787-790. MR1024982 (91f:53035)

\bibitem{r:s2} R. Sharma, {\it Second order parallel tensors on contact manifolds. I},
Algebras Groups Geom., 7(1990), no. 2, 145-152. MR1109567 (92b:53041)

\bibitem{r:s3} R. Sharma, {\it Second order parallel tensors on contact manifolds. II},
C. R. Math. Rep. Acad. Sci. Canada, 13(1991), no. 6, 259-264. MR1145119 (93b:53026)

\bibitem{r:s4} R. Sharma, {\it On the curvature of contact metric manifolds},
J. Geom., 53(1995), no. 1-2, 179-190. MR1337435 (96d:53031)

\bibitem{r:s5} R. Sharma, {\it Certain results on} $K$-{\it contact and} $(k, \mu)$-{\it contact ma\-ni\-folds},
J. Geom., 89(2008), 138-147. MR2457028

\bibitem{h:q} H. Singh; Q. Khan, {\it On spe\-cial weakly symmetric Riemannian manifolds},
Publ. Math. Debrecen, 58(2001), no. 3, 523-536. MR1831059 (2002f:53076)

\bibitem{t:d1} D. Tarafdar; U. C. De, {\it Second order parallel tensors on $P$-Sasakian manifolds},
Northeast. Math. J., 11(1995), no. 3, 260-262. MR1387530 (97c:53075)

\bibitem{t:m} M. Tarafdar; A. Mayra, {\it On nearly Sasakian manifold},
An. Stiint. Univ. "Al. I. Cuza" Ia\c si, 45(1999), no. 2, 291-294. MR1811732 (2001k:53086)

\bibitem{mm:t} M. M. Tripathi, {\it Ricci so\-li\-tons in con\-tact me\-tric ma\-ni\-folds}, \\ arXiv:0801.4222.

\bibitem{y:w} Y.-c. Wong, {\it Existence of linear connections with respect to which
given tensor fields are parallel or recurrent}, Nagoya Math. J., 24(1964), 67-108. MR0174015 (30 \#4222)

\bibitem{h:w} H. Wu, {\it Holonomy groups of indefinite metrics}, Pacific J. Math., 20(1967),
351-392. MR0212740 (35 \#3606)

\bibitem{g:z} G. Zhao, {\it Symmetric covariant tensor fields of order 2 on pseudo-Riemannian manifolds},
Viena Preprint ESI 479 (1997). Available at http://www.esi.ac.at/preprints/esi479.ps
\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{amssymb,amsmath,amsfonts}
\usepackage{color}

\begin{document}

\author{Constantin C\u alin and Mircea Crasmareanu}
\title{From the Eisenhart problem to Ricci solitons in $f$-Kenmotsu manifolds}
\date{Dedicated to the memory of Neculai Papaghiuc 1947-2008}
\maketitle

\begin{abstract}
The Eisenhart problem of finding parallel tensors is solved for the symmetric case in the regular $f$-Kenmotsu framework. On this way, the Olszack-Rosca example of Einstein manifolds provided by $f$-Kenmotsu manifolds via locally symmetric Ricci tensors is recovered as well as a case of Killing vector fields. Some other classes of Einstein-Kenmotsu manifolds are presented. Our result is interpreted in terms of Ricci solitons and special quadratic first integrals.
\end{abstract}

\noindent {\bf 2000 Math. Subject Classification}: 53C40; 53C55; 53C12; 53C42.

\noindent {\bf Key words}: $f$-Kenmotsu manifold; parallel second order covariant tensor field; irreducible metric; Einstein space; Ricci soliton.

\medskip

\section*{Introduction}

In 1923, Eisenhart \cite{l:e} proved that if a positive definite Riemannian ma\-ni\-fold $(M,g)$ admits a second order parallel symmetric covariant tensor other that a constant multiple of the metric tensor, then it is reducible. In 1926, Levy \cite{h:l} proved that a second order parallel symmetric non-degenerated tensor $\alpha $ in a space form is proportional to the metric tensor. Let us point out that this question can be considered as dual to the the problem of finding linear connections making parallel a given tensor field, problem which was considered by Wong in \cite{y:w}. Also, the former question implies topological restrictions namely if the (pseudo) Riemannian manifold $M$ admits a parallel symmetric $(0,2)$ tensor field then $M$ is locally the direct product of a number of (pseudo) Riemannian manifolds, \cite{h:w} (cited by \cite{g:z}). Another situation where the parallelism of $\alpha $ is involved appears in the theory of {\it totally geodesic maps}, namely, as is point out in \cite[p. 114]{o:c}, $\nabla \alpha =0$ is equivalent with the fact that $1:(M,g)\rightarrow (M,\alpha )$ is a totally geodesic map.

While both Eisenhart and Levy work locally, Ramesh Sharma gives in \cite{r:s1} a global approach based on Ricci identities. In addition to space-forms, Sharma considered this {\it Eisenhart problem} in contact geometry \cite{r:s2}-\cite{r:s4}, for example for $K$-contact manifolds in \cite{r:s3}. Since then, several other studies appear in various contact manifolds: nearly Sasakian \cite{t:m}, (para) $P$-Sasakian \cite{t:d1}, \cite{d:e} and \cite{z:l}, $\alpha $-Sasakian \cite{l:d2}. Another framework was that of quasi-constant curvature in \cite{x:j}. Also, contact metrics with nonvanishing $\xi $-sectional curvature are studied in \cite{g:s}.

Returning to contact geometry, an important class of manifolds are introduced by Kenmotsu in \cite{k:k} and generalized by Olszack and Rosca in \cite{o:r}. In the last time, there is an increasing flow of papers in this direction e.g. that of our professor N. Papaghiuc \cite{n:p1}-\cite{n:p2} to which we dedicate this short note. Motivated by this fact we studied the case of $f$-Kenmotsu manifolds satisfying a special condition called by us {\it regular} and show that a symmetric parallel tensor field of second order must be a constant multiple of the Riemannian metric. There are three remarks regarding our result:\\
i) it is in agreement with what happens in all previously recalled contact \\ geometries for the symmetric case, \\
ii) it is obtained in the same manner {\bf as} originated in Sharma's paper \cite{r:s1}, \\
iii) yields a class of Einstein manifolds already indicated by Olszack and Rosca but with a more complicated proof. \\
Let us point out also that the anti-symmetric case appears without proof in \cite{v:m}.

Our main result is connected with the recent theory of Ricci solitons, a subject included in the Hamilton-Perelman approach (and proof)  of Poincar\'e Conjecture. Ricci solitons in contact geometry were first studied by Ramesh Sharma in \cite{g:sc} and \cite{r:s5}; also the preprint \cite{mm:t} is available to arxiv. In these papers the $K$-contact and $(k, \mu )$-contact (including Sasakian) cases are treated; then our treatment for the Kenmotsu variant of almost contact geometry seems to be new.

Our work is structured as follows. The first section is a very brief review of Kenmotsu geometry and Ricci solitons. The next section is devoted to the (symmetric case of) Eisenhart problem in a $f$-Kenmotsu manifold and several situations yielding Einstein manifolds are derived. Also, the relationship with the Ricci solitons is pointed out. The last section offers a dynamical picture of the subject via Killing vector fields and quadratic first integrals of a special type.

{\bf Acknowledgement} Special thanks are offered to Gheorghe Pitis for some useful remarks as well as sending us his book \cite{gh:p}, a source of several references. Also, we are very indebted to Marian-Ioan Munteanu and the referees who pointed out major improvements.

\section{$f$-Kenmotsu manifolds. Ricci solitons}

Let $M$ be a real $2n+1$-dimensional differentiable manifold endowed with an almost
contact metric structure $(\varphi, \xi, \eta, g)$:
$$
\begin{array}{l}
(a) \quad \varphi ^{2} = -I + \eta \otimes \xi,\ \ (b) \quad \eta(\xi) = 1,\ \ (c) \quad \eta \circ \varphi = 0,\\ (d) \quad \varphi (\xi) = 0, \ \
(e) \quad \eta(X) = g(X, \xi),\\  (f) \quad g(\varphi X, \varphi Y)=g(X, Y)-\eta (X)\eta(Y),
\end{array} \eqno(1.1)
$$
for any vector fields $X, Y\in {\cal X}(M)$ where $I$ is the identity of the tangent
bundle $TM$, $\varphi $ is a tensor field of $(1,1)$-type, $\eta $ is a 1-form,
$\xi $ is a vector field and $g$ is a metric tensor field.
Throughout the paper all objects are differentiable of class $C^{\infty }$.

\medskip

We say that $(M,\varphi ,\xi ,\eta ,g)$ is an $f$-{\it Kenmotsu manifold} if the Levi-Civita connection of $g$ satisfy \cite{v:m}:
$$
(\nabla _X\varphi )(Y)=f(g(\varphi X,Y)\xi -\varphi (X)\eta (Y)), \eqno(1.2)
$$
where $f\in C^{\infty }(M)$ is strictly positive and $df\wedge \eta =0$ holds. A $f=constant\equiv \beta >0$ is called $\beta $-{\it Kenmotsu manifold} with the particular case $f\equiv 1$-Kenmotsu manifold which is a usual {\it Kenmotsu manifold} \cite{k:k}.

\medskip

In a general $f$-Kenmotsu manifold we have, \cite{o:r}:
$$
\nabla _X\xi =f(X-\eta (X)\xi), \eqno(1.3)
$$
and the curvature tensor field:
$$
R(X,Y)\xi =f^2(\eta (X)Y-\eta (Y)X)+Y(f)\varphi ^2X-X(f)\varphi ^2Y \eqno(1.4)
$$
while the Ricci curvature and Ricci tensor are, \cite{k:pt}:
$$
S(\xi ,\xi ) = -2n(f^2+\xi (f)) \eqno(1.5)
$$
$$
Q(\xi )=-2nf^2\xi -\xi (f)\xi-(2n-1)grad f. \eqno(1.6)
$$

In the last part of this section we recall the notion of Ricci solitons according to \cite[p. 139]{r:s5}. On the manifold $M$, a {\it Ricci soliton} is a triple $(g,V,\lambda )$ with $g$ a Riemannian metric, $V$ a vector field and $\lambda $ a real scalar such that:
$$
{\cal L}_Vg+2S+2\lambda g=0. \eqno(1.7)
$$
The Ricci soliton is said to be {\it shrinking, steady} or {\it expanding} according as $\lambda $ is negative, zero or positive.

\section{Parallel symmetric second order tensors and \\ Ricci solitons in $f$-Kenmotsu manifolds}

Fix $\alpha $ a symmetric tensor field of $(0,2)$-type which we suppose to be parallel with respect to $\nabla $ i.e. $\nabla \alpha =0$. Applying the Ricci identity \\ $\nabla ^2\alpha (X,Y;Z,W)-\nabla ^2(X,Y;W,Z)=0$ we get the relation $(1.1)$ of \cite[p. 787]{r:s1}:
$$
\alpha (R(X,Y)Z,W) +\alpha (Z,R(X,Y)W)=0, \eqno(2.1)
$$
which is fundamental in all papers treating this subject. Replacing $Z=W=\xi $ and using $(1.4)$ it results:
$$
f^2[\eta (X)\alpha (Y,\xi )-\eta (Y)\alpha (X,\xi)]+Y(f)\alpha
(\varphi ^2X, \xi)-X(f)\alpha (\varphi ^2Y, \xi)=0, \eqno(2.2)
$$
by the symmetry of $\alpha $. With $X=\xi $ we derive:
$$
[f^2+\xi (f)][\alpha (Y,\xi )-\eta (Y)\alpha (\xi , \xi)]=0
$$
and supposing $f^2+\xi (f)\neq 0$ it results:
$$
\alpha (Y,\xi )=\eta (Y)\alpha (\xi , \xi). \eqno(2.3)
$$
Let us call {\it regular} $f$-{\it Kenmotsu manifold} a $f$-Kenmotsu manifold with $f^2+\xi (f)\neq 0$ and remark that $\beta $-Kenmotsu manifolds are regular.

\smallskip

Differentiating the last equation covariantly with respect to $X$ we have:
$$
\alpha (\nabla _XY, \xi )+f[\alpha (X,Y)-\eta (X)\eta (Y)\alpha (\xi ,\xi )]=X(\eta (Y))\alpha (\xi , \xi), \eqno(2.4)
$$
which means via $(2.3)$ with $Y \rightarrow \nabla _XY$:
$$
f[\alpha (X,Y)-\eta (X)\eta (Y)\alpha (\xi ,\xi )]=[X(g(Y,\xi ))-g(\nabla _XY, \xi )]\alpha (\xi ,\xi )=
$$
$$
=g(Y,\nabla _X\xi )\alpha (\xi ,\xi )=
f[g(X,Y)-\eta (X)\eta (Y)]\alpha (\xi ,\xi ). \eqno(2.5)
$$
From the positiveness of $f$ we deduce that:
$$
\alpha (X,Y)=\alpha (\xi ,\xi )g(X,Y) \eqno(2.6)
$$
which together with the standard fact that the parallelism of $\alpha $ implies the constance of $\alpha (\xi ,\xi )$ via $(2.3)$ yields:

\medskip

{\bf Theorem} {\it A symmetric parallel second order covariant tensor in a regular $f$-Kenmotsu manifold is a constant multiple of the metric tensor. In other words, a regular $f$-Kenmotsu metric is irreducible which means that the tangent bundle does not admits a decomposition $TM=E_1\oplus E_2$ parallel with respect of the Levi-Civita connection of $g$}.

\medskip

{\bf Corollary 1} {\it A locally Ricci symmetric} $(\nabla S\equiv 0)$ {\it regular $f$-Kenmotsu manifold is an Einstein manifold}.

\medskip

{\bf Remarks} 1) The particular case of dimension three and $\beta$-Kenmotsu of our theorem appears in Theorem 3.1 from \cite[p. 2689]{d:m}. The above corollary has been proved by Olszack and Rosca in another way. \\
2) In \cite{b:tdt} it is shown the equivalence of the following statements for an Kenmotsu manifold:\\
i) is Einstein, \\
ii) is locally Ricci symmetric, \\
iii) is Ricci semi-symmetric i.e. $R\cdot S=0$ where:
$$
(R(X, Y)\cdot S)(X_1, X_2) = -S(R(X, Y)X_1, X_2)-S(X_1, R(X, Y)X_2).
$$
The same implication iii) $\rightarrow $ i) for Kenmotsu manifolds is Theorem 1 from \cite[p. 438]{j:dp}. But we have the implication iii) $\rightarrow $ i) in the more general framework of regular $f$-Kenmotsu manifols since $R\cdot S=0$ means exactly $(2.1)$ with $\alpha $ replaced by $S$. Every semisymmetric manifold, i. e. $R\cdot R=0$, is Ricci-semisymmetric but the converse statement is not true. In conclusion:

\medskip

{\bf Proposition 1} {\it A Ricci-semisymmetric, particularly semisymmetric, re\-gu\-lar} $f$-{\it Kenmotsu manifold is Einstein.}

\medskip

Another class of spaces related to the Ricci tensor was introduced in \cite{h:q}; namely a Riemannian manifold is a {\it special weakly Ricci symmetric space} if there exists a 1-form $\rho $ such that:
$$
(\nabla _XS)(Y,Z)=2\rho (X)S(Y,Z)+\rho (Y)S(Z,X)+\rho (Z)S(X,Y). \eqno(2.7)
$$
The same condition was sometimes called {\it generalized pseudo-Ricci sym\-metric manifold} (\cite{j:s}) or simply {\it pseudo-Ricci symmetric manifold} (\cite{c:k}). Making $X=Y=Z=\xi $ it results:
$$
\xi (S(\xi ,\xi ))=4\rho (\xi )S(\xi ,\xi) \eqno(2.8)
$$
and then for a $\beta $-Kenmotsu manifold we get $\rho (\xi )=0$. Returning to $(2.7)$ with $Y=Z=\xi $ will results $\rho (X)=0$ for every vector field $X$ and then we have a generalization of Theorem 3.3. from \cite[p. 96]{a:go}:

\medskip

{\bf Proposition 2} {\it A} $\beta $-{\it Kenmotsu manifold which is special weakly Ricci symmetric is an Einstein space.}

\medskip

We close this section with applications of our Theorem to Ricci solitons:

\medskip

{\bf Corollary 2} {\it Suppose that on a regular $f$-Kenmotsu manifold the} $(0,2)$-{\it type field} ${\cal L}_Vg+2S$ {\it is parallel where} $V$ {\it is a given vector field. Then} $(g, V)$ {\it yield a Ricci soliton. In particular, if the given regular} $f$-{\it Kenmotsu manifold is Ricci-semisymmetric or semisymmetric with} ${\cal L}_Vg$ {\it parallel, we have the same conclusion.}

\medskip

Naturally, two situations appear regarding the vector field $V$: $V\in span{\xi }$ and $V\bot \xi$ but the second class seems far too complex to analyse in practice. For this reason it is appropriate to investigate only the case $V=\xi $.

We are interested in expressions for ${\cal L}_{\xi }g+2S$. A straightforward computation gives:
$$
{\cal L}_{\xi }g(X,Y)=2f(g(X,Y)-\eta (X)\eta (Y))=2fg(\varphi X, \varphi Y). \eqno(2.9)
$$

A general expression of $S$ is known by us only for the the 3-dimensional case and $\eta $-Einstein Kenmotsu manifolds. Let us treat these situations in the following:

\medskip

I) \cite[p. 251]{d:t}:
$$
S(X,Y)=\left(\frac{r}{2}+\xi (f)+f^2\right)g(X,Y)-
$$
$$
-\left(\frac{r}{2}+\xi (f)+3f^2\right)\eta (X)\eta (Y)-Y(f)\eta (X)-X(f)\eta (Y) \eqno(2.10)
$$
where $r$ is the scalar curvature. Then, for a 3-dimensional $f$-Kenmotsu manifold we get:
$$
\alpha:=({\cal L}_{\xi }g+2S)(X,Y)=(r+2\xi (f)+2f+2f^2)g(X,Y)-
$$
$$
-(r+2\xi (f)+2f+6f^2)\eta (X)\eta (Y)-2Y(f)\eta (X)-2X(f)\eta (Y) \eqno(2.11)
$$
while, for $\beta $-Kenmotsu:
$$
\alpha (X,Y)=(r+2\beta +2\beta ^2)g(\varphi X, \varphi Y)-4\beta ^2\eta (X)\eta (Y), \eqno(2.12)
$$
$$
(\nabla _Z\alpha )(X,Y)=Z(r)g(\varphi X, \varphi Y)-
$$
$$
-\beta (r+2\beta +6\beta ^2)[\eta (X)g(\varphi Y, \varphi Z)+\eta (Y)g(\varphi X, \varphi Z)]. \eqno(2.13)
$$
{\bf Substituting} $Z=\xi , X=Y\in (span{\xi })^{\bot }$ respectively $X=Y=Z\in (span{\xi })^{\bot }$ in $(2.13)$ we derive that $r$ is a constant, provided $\alpha $ is parallel. Thus, we can state the following:

\medskip

{\bf Proposition 3} {\it A 3-dimensional} ${\beta }$-{\it Kenmotsu Ricci soliton} $(g, \xi , \lambda )$ {\it is expanding and with constant scalar curvature}.

\medskip

{\bf Proof} $\lambda =-\frac{1}{2}\alpha (\xi, \xi)=2\beta ^2$. \quad $\Box $

\medskip

At this point we remark that the Ricci solitons of almost contact geometry studied in \cite{r:s5} and \cite{mm:t} in relationship with the Sasakian case are shrinking and this observation is in accordance with the diagram of Chinea from \cite{d:c} that Sasakian and Kenmotsu are opposite sides of the trans-Sasakian moon. Also, the expanding character may be considered as a manifestation of the fact that a $\beta $-Kenmotsu manifold can not be compact.

\medskip

II) Recall that the metric $g$ is called $\eta $-{\it Einstein} if there exists two real functions $a, b$ such that the Ricci tensor of $g$ is:
$$
S=ag+b\eta \otimes \eta .
$$
For an $\eta $-Einstein Kenmotsu manifold we have, \cite[p. 441]{j:dp}:
$$
S(X,Y)=\left(\frac{r}{2n}+1\right)g(X,Y)-\left(\frac{r}{2n}+2n+1\right)\eta (X)\eta (Y) \eqno(2.14)
$$
and then:
$$
\alpha (X,Y)=\left(\frac{r}{n}+4\right)g(X,Y)-\left(\frac{r}{n}+4+4n\right)\eta (X)\eta (Y) \eqno(2.15)
$$
$$
(\nabla _Z\alpha )(X,Y)=\frac{1}{n}Z(r)g(\varphi X, \varphi Y)-
$$
$$
-\left(\frac{r}{n}+4n+4\right)[\eta (Y)g(\varphi X, \varphi Z)+\eta (X)g(\varphi Y, \varphi Z)]. \eqno(2.16)
$$

\medskip

{\bf Proposition 4} {\it An} ${\eta }$-{\it Einstein Kenmotsu Ricci soliton} $(g, \xi , \lambda )$ {\it is ex\-pan\-ding and with constant scalar curvature, thus Einstein}.

\medskip

{\bf Proof} $\lambda =-\frac{1}{2}\alpha (\xi, \xi)=2n$. The same computation as in Proposition 3 implies constant scalar curvature. \quad $\Box $

\section{The dynamical point of view}

We begin this section with a straightforward consequence of the main Theorem, which also appears in the Olzack-Rosca paper, and is related to the last part of Corollary 2:

\medskip

{\bf Corollary 3} {\it An affine Killing vector field in a $\beta $-Kenmotsu manifold is Killing. As consequence, that scalar provided by the Ricci soliton} $(g, V)$ {\it of a Ricci-semisymmetric} $\beta $-{\it Kenmotsu manifold is} $\lambda =-S(V,V)$.

\medskip

{\bf Proof} (inspired by \cite[p. 504]{g:s}) Fix $X\in {\cal X}(M)$ an affine Killing vector field: $\nabla {\cal L}_Xg=0$. From Theorem it results that $X$ is {\it conformal Killing} i.e. ${\cal L}_Xg = cg$; more precisely $X$ is {\it homothetic} since $c$ is a constant. Lie differentiating the identity
$(1.5)$ along $X$ and using ${\cal L}_XS = 0$ (since $X$ is homothetic) and equation $(1.6)$
we get $g({\cal L}_X\xi ,\xi ) = 0$. Hence $c =({\cal L}_Xg)(\xi ,\xi )=-2g({\cal L}_X\xi ,\xi )= 0$. Thus $X$ is Killing. \quad $\Box $

\medskip

Let us present another dynamical picture of our results. Let $(M, \nabla)$ be a $m$-dimensional manifold endowed with a symmetric linear connection. A {\it quadratic first integral} (QFI on short) for the geodesics of $\nabla $ is defined by ${\cal F}=a_{ij}\frac{dx^i}{dt}\frac{dx^j}{dt}$ with a symmetric 2-tensor field $a=(a_{ij})$ satisfying the {\it Killing-type equations}:
$$
a_{ij:k}+a_{jk:i}+a_{ki:j}=0, \eqno(3.1)
$$
where, as usual, the double dot means the covariant derivative with respect to $\nabla $.

\medskip

The QFI defined by $a$ is called {\it special} (SQFI) if
$a_{ij:k}=0$ and the maximum number of linearly independent SQFI a pair $(M, \nabla)$ can admit
is $\frac{m(m+1)}{2}$; a flat space will admit this number.
In \cite[p. 117]{l:k} it is shown that a non-flat Riemannian
manifold may admit as many as $M_S(m)=1+\frac{(m-2)(m-1)}{2}$
linearly independent SQFI. Therefore, for an almost contact manifold
($m=2n+1$) the maximum number of SQFI is $M_S(2n+1)=1+n(2n-1)>1$.

\medskip

Our main result implies that for a regular $f$-Kenmotsu manifold the number of SQFI is exactly 1 and the only SQFI is the {\it kinetic energy} ${\cal F}=g_{ij}\frac{dx^i}{dt}\frac{dx^j}{dt}$. So:

\medskip

{\bf Proposition 5} {\it There exist almost contact manifolds which does not admit $M_S(2n+1)$ SQFI}.

\medskip

It remains as an open problem to find examples of almost contact metrics with exactly $M_S(2n+1)$ SQFI.

\medskip

\noindent Department of Mathematics, \newline
Technical University "Gh.Asachi"\newline
Ia\c si, 700049\newline
Romania\newline
e-mail: c0nstc@yahoo.com

\medskip

\noindent Faculty of Mathematics
\newline University "Al. I.Cuza" \newline
Ia\c si, 700506 \newline Romania \newline e-mail: mcrasm@uaic.ro

\smallskip

\noindent http://www.math.uaic.ro/$\sim$mcrasm

\end{document}